# Explicit Motion Handling and Interactive Prompting for Video Camouflaged Object Detection

### v1

@@ 3.2 Segmentation-to-Motion Prompt
Position of prompt. The selection of the prompt position is guided by two key observations. Firstly, the choice of the fundamental model plays a pivotal role. Each fundamental model is tailored with a specific architecture; in the case of GMFlow within our EMIP framework, its transformer blocks synergistically form a functional unit crafted for calculating the similarity of pixel features between two frames. Thus, preserving the integrity of this architecture becomes paramount for the coherent generation of optical flow. Secondly, the architectural arrangement of neural networks involves a stratification where lower layers are dedicated to extracting basic features such as edges, colors, and textures, each exhibiting distinctive characteristics across various modalities [?]. This stratification has been shown by [?] to be particularly advantageous for task like optical flow estimation, emphasizing the importance of low-level visual correspondences. Leveraging this insight, we position the segmentation prompt at the lower layers of both the segmentation and motion modeling streams, with the direction being from segmentation to motion. Further details are discussed in Section 4.3, where we delve into various prompt position choices and their impact on results.
Camouflage feeder. To better incorporate the segmentation prompt into the motion stream, we design camouflage feeder. As shown in Fig. 2, our camouflage feeder is based on cross attention with a residual connection, which takes the flow input feature $f^{flow}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of query, and the segmentation prompt feature $f^{2}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of key and value. The residual connection is to maintain more query-related cues, and $d$ represents the embedding dimension. This process can be written as:
@@
and $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ represent query, key, and value matrices, respectively. $\bf Q/\bf K$ and $\bf V$ are derived by applying layer normalization and 3$\times$3 convolution to the input feature maps $f^{flow}_{t}$ and $f^{2}_{t}$, respectively. Here, we employ a Gated-Dconv Feed-forward Network (GDFN) [?] to suppress the aggregation of noise prompt. The GDFN process is defined as:
@@
where $\mathtt{C1}$ and $\mathtt{C3}$ are 1$\times$1 convolution and 3$\times$3 depth-wise convolution, respectively. $\phi$ is GELU function [?] and $\odot$ denotes element-wise multiplication. Note that, $\mathtt{GDFN}$ comprises two separate paths, one of which is activated with GELU function. Subsequently, an element-wise product is applied between these two paths. Literature [?] has demonstrated that the gating mechanism in GDFN can control information flow and yield better performance compared with the conventional feed-forward network (FN) [?] for feature restoration. For the same purpose, to mitigate the impact of the integrated noise prompts, we employ GDFN to control noise information flow. Finally, the resulting segmentation-prompted motion features $\hat{f}^{flow}_{t}$ with both knowledges of motion and appearance can be well adapted to subsequent motion reconstruction.
@@
Prompt learning. To better learn the motion information, inspired by [?], we employ a self-supervised loss to optimize the motion-to-segmentation prompt learning. Further elaboration is provided in Section 3.4.
3.3 Motion-to-Segmentation Prompt
@@
Motion-to-segmentation prompt formally provides an auxiliary motion flow that is both temporally and spatially coordinated with the segmentation stream. After obtaining the segmentation-prompted motion, we need to collect this knowledge back to the segmentation stream. Thus, we introduce another module named Motion Collector to integrate motion prompts into the segmentation stream. Given its functional equivalence of motion collector with the camouflage feeder, we opt to maintain the structural configuration of the interaction block rather than introducing a redesign. Specifically, upon receiving motion feature $\boldsymbol{G}$ (flattening the first two dimensions of $\boldsymbol{M}$), we commence by applying a 3$\times$3 convolution operation to $\boldsymbol{G}$ to adjust the channel dimensions to align with $f^{2}_{t}$. Then we feed the motion-to-segmentation prompt $\hat{\boldsymbol{G}}=\mathtt{Conv}(\boldsymbol{G})$ and $f^{2}_{t}$ to the motion collector to obtain the prompted appearance feature $f_{t\leftarrow t-1}^{2}$.
@@
Subsequently, following that most VCOD works [??] use neighbor connection decoder (NCD) [?] to predict the segmentation map, we use NCD to decode the appearance features. NCD with camouflage feeder can be regarded as the prediction head for the motion fundamental model, similar to the concepts of some previous prompt learning methods [??] that design an extra prediction head after freezing the backbone parameters. Specifically, the appearance feature $f_{t\leftarrow t-1}^{2}$ prompted by motion combined with $f^{3}_{t}$ and $f^{4}_{t}$ are fed to NCD to obtain the predicted map. The optimization of motion-to-segmentation prompt and prediction head is under the constraint of segmentation loss, and more details will be given in Sec. 3.4.
@@
Discussion. The motivation for our prompt learning strategy, freezing the motion stream and fully tuning the segmentation stream, is from the observed performance degradation of the PVT [?] pre-trained on ImageNet when it is directly applied to the VCOD task. Similar challenges are also encountered with large models like SAM [?], which struggle in camouflage scenes [?]. Currently, there lacks a freezable foundational model that excels on the VCOD task. In contrast, we find that the optical flow model could achieve good generalization in most VCOD scenarios, when most of its parameters are frozen.
3.4 Supervision and Loss Function
@@
To ensure optimization for each component of the model, we separately define loss functions for the motion and segmentation stream, and then use these two losses to optimize the entire model jointly. Considering the lack of ground truth (GT) optical flow in video camouflaged scenarios, we propose a self-supervised strategy to learn the optical flow estimation stream. For the sake of brevity and clarity, we refer to [?] and warp the frame $I_{t-1}$ according to the optical flow estimation $\boldsymbol{V}$ to obtain the reconstructed reference frame for $I_{t}$, which is denoted as $\hat{I}_{t}$, i.e., $\hat{I}_{t}=\mathtt{Warp}(\boldsymbol{V},I_{t-1})$. We supervise the optical flow by computing the distance between $I_{t}$ and $\hat{I}_{t}$. Thus, the flow loss $\mathcal{L}_{\text{flow}}$ can be expressed as:
@@
where $\mathtt{SSIM}$ is pixel-wise photometric loss [?], which is commonly used for reflecting image distortion from three aspects: brightness, contrast, and structure.
For the segmentation stream, we employ a hybrid loss function [?], the same as SLT-Net [?]. The hybrid loss $\mathcal{L}_{\text{seg}}$ can be defined as $\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{IoU}}+\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{e-loss}}$, which includes IoU loss, binary cross-entropy loss, and enhanced-alignment loss. Finally, we jointly optimize the motion modeling and object segmentation streams, whose total loss is formulated as:
@@
3.5 Long-term Consistency Modeling
Figure 3: Long-term modeling scheme (${}^{\dagger}$).
@@ 3.5 Long-term Consistency Modeling
Fig. 3 illustrates our long-term consistency modeling scheme, termed ${}^{\dagger}$. Specifically, during training, we freeze the short-term model after it converges and then add the long-term memory module for further learning. Given a video sequence {${{I}_{1},...,{I}_{t-1},{I}_{t}},t>1$}, we sequentially input current frame ${I}_{t}$ and preceding frame ${I}_{t-1}$ into the short-term model to obtain the appearance features $f^{2}_{t}$ and the motion-to-segmentation prompt $\hat{\boldsymbol{G}}$ between the two frames. The element-wise sum of $f^{2}_{t}$ and $\hat{\boldsymbol{G}}$ is conducted, and the results are then fed to an encoder consisting of two convolutional layers. The processed information is later stored into the memory pool. Considering the computational burden and minor impact of distant frames, we implement a FIFO (First-In First-Out) memory pool with fixed capacity ${L}$. Note that such a fixed capacity is configurable according to practical applications. Referring to the long-term setting of SLT-Net, we set ${L}$ to 5 in this paper. Since the memory pool stores appearance and motion features for both the current and historical frames, the features extracted from the current frame are then employed to query the memory pool using a space-time memory read block (STM) [?], yielding the long-term prompt ${P}_{L}$. Next, ${P}_{L}$ and $f^{2}_{t}$ are fed to the motion collector (re-use the same motion collector of the short-term period) for interaction, and the output of collector is subsequently fed to NCD for decoding.
@@
4 Experiments
4.1 Implementation Details and Datasets
For fair comparisons, following the training strategy with [?], we train our model with a two-stage pipeline: first train the backbone on the static training set of COD10K (3,040 images) [?], and then fine-tune the whole model with the temporal components on the training set of MoCA-Mask (19,313 frames) [?]. Evaluation is performed on two widely recognized VCOD benchmarks: CAD (836 frames) [?] and MoCA-Mask test set (3,626 frames). Five evaluation metrics are adopted to assess model performance, i.e., structure measure ($\mathcal{S}_{\alpha}$[?], weighted F-measure ($F_{\beta}^{w}$[?], mean absolute error ($\mathcal{M}$), and mean value of Dice and IoU. The proposed EMIP is implemented by PyTorch [?] on a single NVIDIA 4090 GPU. During the training stage, the batch size is set to 6, and the maximum training epoch is 60. The entire model is optimized using the Adam optimizer [?] with a cosine annealing strategy. The maximum and minimum learning rates, along with the maximum adjusted iterations, are set to 1e-5, 1e-6, and 20, respectively. All the input images are resized to 352$\times$352 and subjected to data augmentation techniques, including color enhancement and random rotation. For inference, EMIP takes any two consecutive frames as input and outputs the segmentation prediction together with the corresponding optical flow estimation. However, for ${}^{\dagger}$, it requires consecutive-frame pairs to be input sequentially, in order to leverage its long-term property.
@@ 4.1 Implementation Details and Datasets
Image-based SINet [?] CVPR’20 0.598 0.231 0.028 0.276 0.202 0.636 0.346 0.041 0.381 0.283
SINet-V2 [?] TPAMI’22 0.588 0.204 0.031 0.245 0.180 0.653 0.382 0.039 0.413 0.318
DGNet [?] MIR’23 0.581 0.184 0.024 0.222 0.156 0.686 0.416 0.037 0.456 0.340
FEDER [?] CVPR’23 0.560 0.165 0.031 0.194 0.137 0.691 0.444 0.029 0.474 0.375
HitNet [?] AAAI’23 0.623 0.299 0.019 0.318 0.254 0.685 0.463 0.031 0.478 0.373
Video-based RCRNet [?] ICCV’19 0.555 0.138 0.033 0.171 0.116 0.627 0.287 0.048 0.309 0.229
MG [?] ICCV’21 0.530 0.168 0.067 0.181 0.127 0.594 0.336 0.059 0.368 0.268
PNS-Net [?] MICCAI’21 0.544 0.097 0.033 0.121 0.101 0.655 0.325 0.048 0.384 0.290
SLT-Net [?] CVPR’22 0.637 0.304 0.027 0.356 0.271 0.696 0.471 0.031 0.484 0.392
${}^{\dagger}$ [?] CVPR’22 0.631 0.311 0.027 0.360 0.272 0.696 0.481 0.030 0.493 0.401
EMIP (Ours) 0.669 0.374 0.017 0.424 0.326 0.710 0.504 0.029 0.528 0.415
${}^{\dagger}$ (Ours) 0.675 0.381 0.015 0.426 0.333 0.719 0.514 0.028 0.536 0.425
Figure 4: Visual comparisons of our models (EMIP and ${}^{\dagger}$) with eight state-of-the-art methods.
4.2 Comparisons with State-of-the-arts
To demonstrate the effectiveness of our method, we compare our model with 10 cutting-edge methods, comprising five image-based and five video-based. To ensure fair evaluation, predictions from these methods are either directly downloaded from the official website or generated using models retrained with open-source code.
Quantitative results are presented in Table 1. Notable performance gains can be observed of EMIP and ${}^{\dagger}$ over recently proposed image-based COD and VCOD models, like HitNet [?] and SLT-Net [?], across all five evaluation metrics. Several visual samples are shown in Fig. 4. EMIP appears to be more effective at simultaneously conducting motion and segmentation, and incorporating both enhanced features further makes it better for detecting camouflaged objects. Additionally, predictions of consecutive frames are listed in Fig. 6. One can see that incorporating the long-term modeling scheme, namely ${}^{\dagger}$, can further refine results and lead to boosted performance.

### v2

@@ III-B Segmentation-to-Motion Prompt
III-B1 Position of prompt
The selection of the prompt position is guided by two key observations. Firstly, the choice of the fundamental model plays a pivotal role. Each fundamental model is tailored with a specific architecture; in the case of GMFlow within our EMIP framework, its transformer blocks synergistically form a functional unit crafted for calculating the similarity of pixel features between two frames. Thus, preserving the integrity of this architecture becomes paramount for the coherent generation of optical flow. Secondly, the architectural arrangement of neural networks involves a stratification where lower layers are dedicated to extracting basic features such as edges, colors, and textures, each exhibiting distinctive characteristics across various modalities [62]. This stratification has been shown by [15] to be particularly advantageous for task like optical flow estimation, emphasizing the importance of low-level visual correspondences. Leveraging this insight, we position the segmentation prompt at the lower layers of both the segmentation and motion modeling streams, with the direction being from segmentation to motion. Further details are discussed in Section IV-D, where we delve into various prompt position choices and their impact on results.
III-B2 Camouflage feeder
To better incorporate the segmentation prompt into the motion stream, we design camouflage feeder. As shown in Fig. 2 right, our camouflage feeder is based on cross attention with a residual connection, which takes the optical flow input feature $f^{flow}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of query, and the segmentation prompt feature $f^{2}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of key and value. The residual connection is to maintain more query-related cues, and $d$ represents the embedding dimension. This process can be written as:
@@ III-B2 Camouflage feeder
and $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ represent query, key, and value matrices, respectively. $\bf Q/\bf K$ and $\bf V$ are derived by applying layer normalization and 3$\times$3 convolution to the input $f^{flow}_{t}$ and $f^{2}_{t}$, respectively. Here, we employ a Gated-Dconv Feed-forward Network (GDFN) [63] to suppress the aggregation of noisy prompt, and GDFN for input $\mathbf{X}$ is defined as:
@@
where $\mathtt{C1}$ and $\mathtt{C3}$ are 1$\times$1 convolution and 3$\times$3 depth-wise convolution, respectively. $\phi$ is GELU function [64] and $\odot$ denotes element-wise multiplication. Note that, $\mathtt{GDFN}$ comprises two separate paths, one of which is activated with GELU function. Subsequently, an element-wise product is applied between these two paths. Literature [63] has demonstrated that the gating mechanism in GDFN can control information flow and yield better performance compared with the conventional feed-forward network (FN) [65] for feature restoration. For the same purpose, to mitigate the impact of the integrated noisy prompts, we employ GDFN to control noisy information flow. Finally, the resulting segmentation-prompted motion features $\hat{f}^{flow}_{t}$ with both knowledge of motion and appearance can be well adapted to subsequent motion reconstruction.
III-B3 Prompt learning
@@
To better learn the motion information, inspired by [66], we employ a self-supervised loss to optimize the motion-to-segmentation prompt learning. Further elaboration is provided in Section III-D.
III-C Motion-to-Segmentation Prompt
@@
Motion-to-segmentation prompt formally provides an auxiliary motion flow that is both temporally and spatially coordinated with the segmentation stream. After obtaining the segmentation-prompted motion, we need to collect this knowledge back to the segmentation stream. Thus, we introduce another module named Motion Collector to integrate motion prompts into the segmentation stream. Given the functional analogy of motion collector with the camouflage feeder, we opt to maintain the structural configuration of the interaction block rather than introducing a redesign (as illustrated in Fig. 2 right). Specifically, upon receiving motion feature $\boldsymbol{G}$ (flattening the first two dimensions of $\boldsymbol{M}$), we then apply a 3$\times$3 convolution operation to $\boldsymbol{G}$ to adjust and reduce its channel dimension to align with $f^{2}_{t}$. Then we feed the motion-to-segmentation prompt $\hat{\boldsymbol{G}}=\mathtt{Conv}(\boldsymbol{G})$ and $f^{2}_{t}$ to the motion collector to obtain the prompted appearance feature $f_{t\leftarrow t-1}^{2}$.
@@
Subsequently, following that most VCOD works [9, 10] use neighbor connection decoder (NCD) [10] to predict the segmentation map, we use NCD to decode the appearance features. NCD with motion collector can be regarded as the prediction head for the motion fundamental model, similar to the concepts of some previous prompt learning methods [9, 10] that design an extra prediction head after freezing the backbone parameters. Specifically, the appearance feature $f_{t\leftarrow t-1}^{2}$ prompted by motion combined with $f^{3}_{t}$ and $f^{4}_{t}$ are fed to NCD to obtain the predicted map. The optimization of motion-to-segmentation prompt and prediction head is under the constraint of segmentation loss, and more details will be given in Sec. III-D.
@@
Discussion. The motivation for our prompt learning strategy, freezing the motion stream and fully tuning the segmentation stream, is from the observed performance degradation of the PVT [60] pre-trained on ImageNet when directly applied to the VCOD task. Similar challenges are also encountered with large models like SAM [59], which struggle in camouflaged scenes [67]. Currently, there lacks a freezable foundational model that excels on the VCOD task. In contrast, we find that the optical flow model could achieve good generalization in most VCOD scenarios, when most of its parameters are frozen.
III-D Supervision and Loss Function
@@
To ensure optimization for each component of the model, we separately define loss functions for the motion and segmentation stream, and then use these two losses to optimize the entire model jointly. Considering the lack of ground truth (GT) optical flow in video camouflaged scenarios, we propose a self-supervised strategy to learn the optical flow estimation stream. For the sake of brevity and clarity, we refer to [66] and warp the frame $I_{t-1}$ according to the optical flow estimation $\boldsymbol{V}$ to obtain the reconstructed reference frame for $I_{t}$, which is denoted as $\hat{I}_{t}$, i.e., $\hat{I}_{t}=\mathtt{Warp}(\boldsymbol{V},I_{t-1})$. We supervise the optical flow by computing the distance between $I_{t}$ and $\hat{I}_{t}$. Thus, the flow loss $\mathcal{L}_{\text{flow}}$ can be expressed as:
@@
where $\mathtt{SSIM}$ is pixel-wise photometric loss [68], which is commonly used for reflecting image distortion from three aspects: brightness, contrast, and structure.
For the segmentation stream, we employ a hybrid loss function [69], the same as SLT-Net [10]. The hybrid loss $\mathcal{L}_{\text{seg}}$ is defined as $\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{IoU}}+\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{e-loss}}$, which includes IoU loss, binary cross-entropy loss, and enhanced-alignment loss. Finally, we jointly optimize the motion modeling and object segmentation streams, and the total loss is formulated as:
@@
III-E Long-term Consistency Modeling
Figure 3: Overview of our long-term modeling scheme (EMIP). EMIP consists of a frozen EMIP and other five learning modules (i.e., Memory Encoder, Query Encoder, STM, Motion Collector, and NCD).
@@ III-E Long-term Consistency Modeling
Fig. 3 illustrates our long-term consistency modeling scheme, termed EMIP. Specifically, during training, we freeze the short-term model (referring to the explicit motion modeling stream and the object segmentation stream in Fig. 3) after it converges and then add the long-term memory module for further learning. Given a video sequence {${{I}_{1},...,{I}_{t-1},{I}_{t}},t>1$}, we sequentially input current frame ${I}_{t}$ and preceding frame ${I}_{t-1}$ into the short-term model to obtain the appearance features $f^{2}_{t}$ and the motion-to-segmentation prompt $\hat{\boldsymbol{G}}$ between the two frames. The element-wise sum of $f^{2}_{t}$ and $\hat{\boldsymbol{G}}$ is conducted, and the results are then fed to the memory encoder to be mapped as key $K_{Mt}$ and value $V_{Mt}$. The memory encoder consists of a convolutional layer, followed by a normalization layer and a ReLU activation function, and then two parallel convolutional layers. The input features are processed to obtain the key-value mappings $K_{Mt}$ and $V_{Mt}$. The processed key-value mappings are later stored into the memory pool.
Considering the computational burden and minor impact of distant frames, we implement a FIFO (First-In First-Out) memory pool with fixed capacity ${L}$. Note that such a fixed capacity is configurable according to practical applications. Referring to the long-term setting of SLT-Net [10], we set ${L}$ to 5 in this paper. The object feature $f^{2}_{t}$ extracted from the current frame are fed to the query encoder, which consists of two parallel convolutional layers used to compute the query key-value mappings $K_{Q}$ and $V_{Q}$. These mappings are then used to query the memory pool, which stores appearance and motion features for both the current and historical frames, via a space-time memory read block (STM) [46] (referring to the right part of Fig. 3). This process yields the long-term prompt ${P}_{L}$. Specifically, the key-value mappings of each historical frame in the memory pool are concatenated along the temporal dimension to form $K_{M}$ and $V_{M}$. These concatenated mappings are then interacted with the query key-value mappings $K_{Q}$ and $V_{Q}$ through the STM module, ultimately producing the query results. Next we apply a $3\times 3$ convolution operation to ${P}_{L}$ to adjust and reduce its channel dimension to align with $f^{2}_{t}$. Then, $\hat{P}_{L}=\mathtt{Conv}({P}_{L})$ and $f^{2}_{t}$ are fed to the motion collector [^1] Can re-use the same motion collector of the short-term period if the short-term output is no longer needed. for interaction, and the output of collector, denoted as $f_{t\leftarrow t-L}$, is subsequently fed to NCD for decoding.
@@
IV Experiments and Results
IV-A Datasets and Metrics
IV-A1 Datasets
Following [10, 40], we conduct experiments on two widely recognized VCOD benchmarks: MoCA-Mask [10] and CAD [38]. Among them, MoCA-Mask stands as the most challenging dataset, comprising 19,313 frames across 71 clips for training, and 3,626 frames of 16 clips for testing. The CAD dataset includes 836 frames of 9 clips designated for testing. To assess the generalizability of our model, we conducted experiments on four widely used VSOD/VOS datasets: DAVIS16 [70] (with 30 training clips and 20 testing clips), FBMS [71] (29 training clips and 30 testing clips), ViSal [72] (comprising 17 video sequences for testing) and SegV2 [73] (including 13 clips for testing).
IV-A2 Evaluation Metrics
We adopt widely recognized evaluation metrics to assess our model performance, namely: structure measure ($\mathcal{S}_{\alpha}$[74], weighted F-measure ($F_{\beta}^{w}$[75], F-measure (max $\mathcal{F}$[76], mean absolute error ($\mathcal{M}$[77], and mean value of Dice and IoU. These metrics provide a comprehensive and reliable assessment of model performance.
IV-B Implementation Details
For fair comparisons, we adhere to the training settings outlined in [10] and employ PVTv2 [60] as the feature extraction backbone. All input images are resized to 352$\times$352 and subject to data augmentation techniques, including color enhancement and random rotation. A two-stage training pipeline in [10] is also adopted by our model: first train the backbone on the static training set of COD10K (3,040 images) [7], and then fine-tune the whole model with the temporal components on the training set of MoCA-Mask (19,313 frames) [10]. The entire model is optimized using the Adam optimizer [78] with a cosine annealing strategy. The maximum and minimum learning rates, along with the maximum adjusted iterations, are set to 1e-5, 1e-6, and 20, respectively. EMIP is trained for 7 hours over 60 epochs on an NVIDIA 4090 GPU with 16,536 MB of memory using a batch size of 6. For inference, EMIP takes any two consecutive frames as input and outputs the segmentation prediction together with the corresponding optical flow estimation. However, for EMIP, it requires consecutive-frame pairs to be input sequentially, in order to leverage its long-term property. For VSOD, we train our model on the training set of DAVIS16 (30 clips) and FBMS (29 clips) as [56]. The proposed EMIP is implemented by PyTorch [79], and all experiments are conducted on an NVIDIA RTX 4090 GPU.
@@ IV-B Implementation Details
Image-based EGNet [80] ICCV’19 VGG-16 0.547 0.110 0.035 0.143 0.096 0.619 0.298 0.044 0.324 0.243
BASNet [81] CVPR’19 ResNet-34 0.561 0.154 0.042 0.190 0.137 0.639 0.349 0.054 0.393 0.293
CPD [82] CVPR’19 VGG-16 0.561 0.121 0.041 0.162 0.113 0.622 0.289 0.049 0.330 0.239
PraNet [61] MICCAI’20 Res2Net-50 0.614 0.266 0.030 0.311 0.234 0.629 0.352 0.042 0.378 0.290
SINet [8] CVPR’20 ResNet-50 0.598 0.231 0.028 0.276 0.202 0.636 0.346 0.041 0.381 0.283
SINet-V2 [7] TPAMI’22 ResNet-50 0.588 0.204 0.031 0.245 0.180 0.653 0.382 0.039 0.413 0.318
ZoomNet [28] CVPR’22 ResNet-50 0.582 0.211 0.033 0.224 0.167 0.587 0.225 0.063 0.246 0.166
DGNet [9] MIR’23 PVT 0.581 0.184 0.024 0.222 0.156 0.686 0.416 0.037 0.456 0.340
FEDER [34] CVPR’23 ResNet-50 0.560 0.165 0.031 0.194 0.137 0.691 0.444 0.029 0.474 0.375
FSPNet [31] CVPR’23 ViT 0.594 0.182 0.044 0.238 0.167 0.539 0.220 0.145 0.309 0.212
PUENet [32] TIP’23 ViT 0.594 0.204 0.037 0.302 0.212 0.673 0.427 0.034 0.499 0.389
HitNet [33] AAAI’23 PVT 0.623 0.299 0.019 0.318 0.254 0.685 0.463 0.031 0.478 0.373
FSEL [83] ECCV’24 PVT 0.596 0.260 0.053 0.219 0.151 0.649 0.368 0.053 0.434 0.325
HGINet [36] TIP’24 ViT 0.610 0.251 0.030 0.303 0.221 0.680 0.437 0.050 0.501 0.392
Video-based RCRNet [52] ICCV’19 ResNet-50 0.555 0.138 0.033 0.171 0.116 0.627 0.287 0.048 0.309 0.229
MG [13] ICCV’21 VGG-style 0.530 0.168 0.067 0.181 0.127 0.594 0.336 0.059 0.368 0.268
PNS-Net [84] MICCAI’21 Res2Net-50 0.544 0.097 0.033 0.121 0.101 0.655 0.325 0.048 0.384 0.290
SLT-Net [10] CVPR’22 PVT 0.637 0.304 0.027 0.356 0.271 0.696 0.471 0.031 0.484 0.392
SLT-Net [10] CVPR’22 PVT 0.631 0.311 0.027 0.360 0.272 0.696 0.481 0.030 0.493 0.401
IMEX [40] TMM’24 ResNet-50 0.661 0.371 0.020 0.409 0.319 0.684 0.452 0.033 0.469 0.370
EMIP (Ours) PVT 0.669 0.374 0.017 0.424 0.326 0.710 0.504 0.029 0.528 0.415
EMIP (Ours) PVT 0.675 0.381 0.015 0.426 0.333 0.719 0.514 0.028 0.536 0.425
Figure 4: Visual comparisons of our models (EMIP and EMIP) with eight state-of-the-art methods. We select some difficult scenarios, including dusky night, fast-moving objects, stationary objects, small objects, and noisy backgrounds.
IV-C Comparisons with State-of-the-arts
To demonstrate the effectiveness of our EMIP, we compare it with various state-of-the-arts. These compared methods can be categorized into two types: (i) Image-based camouflaged object detection methods [80, 81, 82, 61, 8, 7, 28, 9, 34, 31, 32, 33], which are designed for detecting objects in static camouflaged scenes; (ii) Video-based object detection methods [52, 13, 84, 10, 40], which focus on identifying camouflaged or moving objects in dynamic video sequences. Notably, we include comparisons with SLT-Net [10] and IMEX [40], two most relevant state-of-the-art approaches for video camouflaged object detection. For fair evaluation, the predictions from these methods are either directly downloaded from the official repositories or generated using open-source code provided by the original authors.
IV-C1 Quantitative Evaluation
Table I presents the experimental results on VCOD datasets. Notably, on MoCA-Mask dataset, our method demonstrates significant improvements: (i) It surpasses the previous state-of-the-art image-based approach, HitNet [33], by 27% in $F_{\beta}^{w}$, underscoring the efficacy of incorporating motion cues into the VCOD task. (ii) Additionally, when compared to SLT-Net [10], which utilizes the same backbone as ours, our method shows a marked improvement, achieving a 20% gain in $F_{\beta}^{w}$. Although our model was trained on the MoCA-Mask dataset, the testing results on the CAD dataset underscore its superior robustness and quantitative performance enhancements. The notable performance gains of EMIP over existing and recently proposed techniques emphasize that the combination of explicit motion modeling and prompt learning substantially enhances the completeness of detected camouflaged objects.
Additionally, by leveraging long-term temporal information, our model achieves state-of-the-art performance across all five evaluation metrics on both VCOD benchmarks. This indicates that preserving long-term consistency effectively suppresses motion noise of camouflaged objects and enhances the stability of video predictions. Compared to SLT-Net [10] and IMEX [40], which also utilizes long-term consistency information as external cues to refine predictions, our long-term strategy yields notable performance gains. It is worth noting that, unlike SLT-Net, which frames long-term modeling as a sequence-to-sequence problem (requiring all frames of a clip), our approach utilizes only historical frames relative to the current frame. This design makes our scheme theoretically more suitable for real-time and practical applications, where only past information is accessible.
IV-C2 Qualitative Comparison
Fig. 4 illustrates qualitative comparisons by visualizing the segmentation results of several examples. Our model demonstrates superior alignment with ground truth, showcasing its enhanced capability in identifying camouflaged objects compared to other methods. Furthermore, the segmentation results over consecutive frames of the same clip are shown in Fig. 5. One can see that incorporating the long-term modeling scheme, namely EMIP, can reduce errors in short-term prediction and lead to boosted performance. Fig. 11 presents visual comparisons of optical flow prediction in camouflaged scenarios. The designed prompting strategy in our model ensures more precise responses within camouflaged regions, while minimizing the influence of irrelevant motion noise. Consequently, our model excels in concurrently performing the motion modeling and segmentation tasks, leveraging the integration of enhanced features to improve the detection of camouflaged objects.
Figure 5: Visual comparisons on consecutive video frames.
IV-D Ablation Studies
To evaluate the effectiveness of each core component, we conduct thorough ablation studies by removing or substituting components from the complete EMIP.
IV-D1 Camouflage feeder and motion collector
As shown in Table II, we first validate the role of the two prompt integration modules, i.e., camouflage feeder and motion collector. We start with the baseline model (#1), which directly uses the segmentation stream to extract single-frame information for prediction, i.e., removing the motion modeling stream of EMIP. Configuration #2 means using only motion information to prompt the segmentation stream, meanwhile discarding segmentation-to-motion prompt. Model #3 is the full EMIP. Comparing configuration #2 with #1, it is evident that camouflaged objects are challenging to detect without inter-frame motion information. In scenes with movement, motion information provides additional context that can improve the accuracy of segmentation. It helps in identifying and separating camouflaged objects that are in motion, which might be challenging using appearance-based methods alone. Comparing configuration #3 with #2, we observe that incorporating the segmentation-to-motion prompt via the camouflage feeder into the motion stream enhances the robustness of motion cues. This improvement in motion cue robustness subsequently leads to superior performance in camouflaged object detection. The visualization results are presented in Fig. 6. Relying solely on appearance features makes it challenging to accurately detect or localize truly camouflaged objects in certain scenarios. With the introduction of the motion collector, the camouflaged object is detected using motion information. Furthermore, incorporating the camouflage feeder to inject camouflage priors effectively reduces interference and enables more robust localization of the camouflaged object.